OpenStudy (aivantettet26):

A 3,500 kg truck accelerates uniformly from 15.00 m/s to 35.00 m/s east. During this acceleration, the track travels 187 m. What is the net force acting on the car during this acceleration.

OpenStudy (maths99):

Calculate the trucks acceleration then plug that into F=ma.

OpenStudy (g3n3sis08):

Correct me if I'm wrong. a=(v-u)/t 35-15/t = 20/t s=ut + 1/2 at^2 187=(15)t+(0.5x(20/t)xt^2) 187=15t + 10t 187/25=t a=20/t ... (20x25)/187=a F=ma F=(3500x20x25)/187 F=9358.3 N therefore, F is approx. 9360 N